OpenStudy (anonymous):

A cuboidal room is 1 and 1/2 as long again as it is broad. The cost of carpeting the room at the rate of Rs.5 per m sq is Rs.480 and the cost of papering the walls with coloured paper at the rate of Rs.3 per m sq is Rs.360. Find the dimensions of tne room .

OpenStudy (anonymous):

Let breadth of the room be b Let length of the room be l and let height of the roo be h then, Area of floor =l*b 480/5 =l*b l*b=96 ....i Area of 4 walls = 2*(l*h) +2*(b*h) 360/3 =2h(l+b) h(l+b)=60 ....ii l=(1+1/2)b ....iii

OpenStudy (anonymous):

U need to solve that three equations

OpenStudy (anonymous):

l = 12 m , b = 8m, h = 3m